# Diff-Instruct++: Training One-step Text-to-image Generator Model to Align with Human Preferences

### v1

@@ Abstract
One-step text-to-image generator models offer advantages such as swift inference efficiency, flexible architectures, and state-of-the-art generation performance. In this paper, we study the problem of aligning one-step generator models with human preferences for the first time. Inspired by the success of reinforcement learning using human feedback (RLHF), we formulate the alignment problem as maximizing expected human reward functions while adding an Integral Kullback-Leibler divergence term to prevent the generator from diverging. By overcoming technical challenges, we introduce Diff-Instruct++ (DI++), the first, fast-converging and image data-free human preference alignment method for one-step text-to-image generators. We also introduce novel theoretical insights, showing that using CFG for diffusion distillation is secretly doing RLHF with DI++. Such an interesting finding brings understanding and potential contributions to future research involving CFG. In the experiment sections, we align both UNet-based and DiT-based one-step generators using DI++, which use the Stable Diffusion 1.5 and the PixelArt-$\alpha$ as the reference diffusion processes. The resulting DiT-based one-step text-to-image model achieves a strong Aesthetic Score of 6.19 and an Image Reward of 1.24 on the COCO validation prompt dataset. It also achieves a leading Human preference Score (HPSv2.0) of 28.48, outperforming other open-sourced models such as Stable Diffusion XL, DMD2, SD-Turbo, as well as PixelArt-$\alpha$. Both theoretical contributions and empirical evidence indicate that DI++ is a strong human-preference alignment approach for one-step text-to-image models.
@@ 1 Introductions
In recent years, deep generative models have achieved remarkable successes across various data generation and manipulation applications (Karras et al., 2020; 2022; Nichol & Dhariwal, 2021; Oord et al., 2016; Ho et al., 2022; Poole et al., 2022; Hoogeboom et al., 2022; Kim et al., 2022; Tashiro et al., 2021; Kingma & Dhariwal, 2018; Chen et al., 2019; Meng et al., 2021; Couairon et al., 2022; Zhang et al., 2023a; Luo & Zhang, 2024; Xue et al., 2023; Luo et al., 2023b; Zhang et al., 2023b; Feng et al., 2023; Deng et al., 2024; Luo et al., 2024c; Geng et al., 2024; Wang et al., 2024; Pokle et al., 2022). These models have notably excelled in producing high-resolution, text-conditional models such as images (Rombach et al., 2022; Saharia et al., 2022; Ramesh et al., 2022; 2021; Luo et al., 2024b) and other modalities with different applications(Brooks et al., 2024; Kondratyuk et al., 2023; Evans et al., 2024; Luo et al., 2024c), pushing the boundaries of Artificial Intelligence Generated Content.
@@ 3.3 A Practical Algorithm
while not converge do
       fix $\theta$, update ${\psi}$ for $K_{TA}$ rounds by
  1. 1.
  2. 2.
  3. 3.
fix $\psi$, update $\theta$ using StaD:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
end while
return $\theta,\psi$.
Algorithm 2 Diff-Instruct++ Pseudo Code.
@@ 5.3 Qualitative Evaluations
Human Reward Improves the One-step Model.
Figure 3 shows a qualitative comparison of five one-step generator models with different alignment configurations. The bottom row of Figure 3 is the weakest setting with no human preference alignment. Upper rows are models that are aligned stronger in a progressive way. From Figure 3 we can tell some findings: (1) The model with no human preference alignment (the bottom row) shows weak performances. Though it can generate acceptable images of scenes, it can barely generate high-quality humans or animals; (2) Without CFG, the model aligned with only human reward model is pretty good, with a significant improvement over the model with no human preference alignment; (3) The model aligned only with CFG is a decent solution; (4) The model aligned with both CFG and weak human reward model shows better image composition, richer details, and aesthetic quality; (5) The model aligned with CFG and strong human reward tends to generate images similar to paintings that are more colorful with richer details; Due to page limitations, we put a discussion on findings in Appendix A.3.
5.4 Limitations of Diff-Instruct++
Despite the alignment training stage, the generator model still makes simple mistakes occasionally. Figure 5 shows some bad generation cases that we picked with multiple generation trails. (1) The Aligned Model Still Misunderstands the Input Prompt. As the leftmost three images of Figure 5 show, the generated images ignore the concept of pear earrings, the battling in a coffee cup, and the car playing football. However, we find such a mistake happens only occasionally. (2) The Generator Sometimes Generates Bad Human Faces and Hands. Please see the fourth and fifth images of Figure 5. The face of the generated lady in the fourth image is not satisfying with blurred eyes and mouth. In the fifth image, the generated Iron Man character has multiple hands. (3) Sometimes The aligned model Still Can not Count Correctly. For instance, in the rightmost image, the prompt asks the model to generate a birthday cake, however, the model generates two cakes with one near and another lying far away.
@@ Appendix A Important Materials for Main Content
A.1 Meanings of Hyper-parameters.
@@
A.2 Experiment Details for Pre-training and Alignment
@@
A.3 More Discussions on Findings of Qualitative Evaluations

### v2

@@ Abstract
One-step text-to-image generator models offer advantages such as swift inference efficiency, flexible architectures, and state-of-the-art generation performance. In this paper, we study the problem of aligning one-step generator models with human preferences for the first time. Inspired by the success of reinforcement learning using human feedback (RLHF), we formulate the alignment problem as maximizing expected human reward functions while adding an Integral Kullback-Leibler divergence term to prevent the generator from diverging. By overcoming technical challenges, we introduce Diff-Instruct++ (DI++), the first, fast-converging and image data-free human preference alignment method for one-step text-to-image generators. We also introduce novel theoretical insights, showing that using CFG for diffusion distillation is secretly doing RLHF with DI++. Such an interesting finding brings understanding and potential contributions to future research involving CFG. In the experiment sections, we align both UNet-based and DiT-based one-step generators using DI++, which use the Stable Diffusion 1.5 and the PixelArt-$\alpha$ as the reference diffusion processes. The resulting DiT-based one-step text-to-image model achieves a strong Aesthetic Score of 6.19 and an Image Reward of 1.24 on the COCO validation prompt dataset. It also achieves a leading Human preference Score (HPSv2.0) of 28.48, outperforming other open-sourced models such as Stable Diffusion XL, DMD2, SD-Turbo, as well as PixelArt-$\alpha$. Both theoretical contributions and empirical evidence indicate that DI++ is a strong human-preference alignment approach for one-step text-to-image models. The homepage of the paper is: https://github.com/pkulwj1994/diff_instruct_pp.
@@ 1 Introductions
In recent years, deep generative models have achieved remarkable successes across various data generation and manipulation applications (Karras et al., 2020; 2022; Nichol & Dhariwal, 2021; Oord et al., 2016; Ho et al., 2022; Poole et al., 2022; Hoogeboom et al., 2022; Kim et al., 2022; Tashiro et al., 2021; Kingma & Dhariwal, 2018; Chen et al., 2019; Meng et al., 2021; Couairon et al., 2022; Zhang et al., 2023a; Luo & Zhang, 2024; Xue et al., 2023; Luo et al., 2023c; Zhang et al., 2023b; Feng et al., 2023; Deng et al., 2024; Luo et al., 2024c; Geng et al., 2024; Wang et al., 2024; Pokle et al., 2022). These models have notably excelled in producing high-resolution, text-conditional models such as images (Rombach et al., 2022; Saharia et al., 2022; Ramesh et al., 2022; 2021; Luo et al., 2024b) and other modalities with different applications(Brooks et al., 2024; Kondratyuk et al., 2023; Evans et al., 2024; Luo et al., 2024c), pushing the boundaries of Artificial Intelligence Generated Content.
@@ 5.2 Quantitative Evaluations with Standard Scores
Ablation Comparison
Table 2 shows an ablation comparison of the effects of Diff-Instruct++ with different explicit reward scales $\alpha_{r}$ and classifier-free guidance reward scales $\alpha_{c}$. For instance, for the SD1.5 model, we find that a larger explicit reward scale always leads to higher Image Reward and Aesthetic Score, while it may trade off the CLIPScore which represents the image-text semantic alignment. Such an observation indicates a trade-off between the image-text alignment and human-preference alignment if we solely use human-preference reward as the implicit reward. One possible solution is to also add the image-text alignment scores such as the CLIPScore as an explicit reward. Besides, we also find that a CFG reward scale that is too large might hurt the human-preference alignment. For instance, the SD1.5-DI++ one-step model with $(\alpha_{r}=100,\alpha_{c}=1.5)$ shows better score than models with $(\alpha_{r}=100,\alpha_{c}=4.5)$. This indicates that though CFG is practically useful in diffusion sampling, it is not perfectly aligned with human preferences. Therefore, finding a proper CFG reward scale is important for optimal human-preference alignment with Diff-Instruct++.
@@ 5.3 Qualitative Evaluations
The Effects of Human Reward and CFG Reward.
Figure 3 shows a visualization comparison of DiT-based one-step models trained with DI++ with different image reward scales and CFG reward scales. We will give some intuition to better understand the effects of two rewards.
  1. 1.
    From the bottom to the top, the first row of Figure 3* is the generated results from models trained with no image reward and no CFG reward (i.e., scales are set to 0). These images are pretty realistic (such as the coast image), but are of poor details and aesthetic appearance;
  2. 2.
    The second row is the model trained only with a unit scale of image reward. These images are much better than DI++ models with both rewards. However, we can see that solely using image rewards results in vivid colors. This is possible because of the fact that human beings would prefer colorful images with great details. However, only using image rewards results in the objects in the generated images being small, which is possibly not wanted;
  3. 3.
    The third row is the generated images from models trained only with CFG reward. We can see the resulting images are pretty aesthetic. However, these images have too big objects and worse background details;
  4. 4.
    The fourth row is the generated images from models trained with unit image reward together with CFG reward. These images show improved image layout, vivid colors, as well as rich details in both foreground and background. We think these models are of the best results. The quantitative HPS score in Table 1 also confirms our observations;
  5. 5.
    The fifth row is the images from models trained with large image rewards and standard CFG rewards. We find that these images tend to be more like paintings instead of realistic photos. The colors are also much brighter than other models. This shows that a too-large image reward scale may lead to the model generating painting-like images instead of realistic photo-like images. Therefore, researchers are encouraged to carefully choose the image reward scales and the CFG scales to train one-step models with Diff-Instruct++ for their own purposes.
Figure 1 to Figure 6 in the supplementary material shows more comprehensive qualitative results which may bring a deeper understanding of DI++. Due to page limitations, we put more discussions on qualitative findings in Appendix A.4.
5.4 Limitations and Failure Cases of Diff-Instruct++
Despite the alignment training stage, the generator model still makes simple mistakes occasionally. Figure 5 shows some bad generation cases that we picked with multiple generation trails. (1) The Aligned Model Still Misunderstands the Input Prompt. As the leftmost three images of Figure 5 show, the generated images ignore the concept of pear earrings, the battling in a coffee cup, and the car playing football. However, we find such a mistake happens only occasionally. (2) The Generator Sometimes Generates Bad Human Faces and Hands. Please see the fourth and fifth images of Figure 5. The face of the generated lady in the fourth image is not satisfying with blurred eyes and mouth. In the fifth image, the generated Iron Man character has multiple hands. (3) Sometimes The aligned model Still Can not Count Correctly. For instance, in the rightmost image, the prompt asks the model to generate a birthday cake, however, the model generates two cakes with one near and another lying far away. Besides, as Figure 3 shows, very strong explicit reward scales lead the image to be very colorful with vivid details. This may cause generated images to look more like paintings instead of realistic photos. Therefore, we suggest researchers carefully choose the explicit reward scales according to their alignment purpose when using Diff-Instruct++.
@@ 6 Conclusion and Future Works
Acknowledgments
First, we would like to acknowledge Zhengyang Geng for helpful discussions on experiment settings and the overall writing of the paper. We also appreciate the authors of Diff-Instruct(Luo et al., 2023b), Score-implicit Matching(Luo et al., 2024b) and Long-short Score-identity Distillation(Zhou et al., 2024a) for their high-quality codebases on diffusion distillation. Second, we would also like to acknowledge Xiaohongshu Inc. for its support of computing on ablation experiments. Finally, we would like to thank the reviewers and editors of the Diff-Instruct++ paper for their professional opinions and wonderful efforts in organizing the reviewing process.
@@ Appendix A Important Materials for Main Content
A.1 Pseudo-code of Diff-Instruct++
while not converge do
       fix $\theta$, update ${\psi}$ for $K_{TA}$ rounds by
  1. 1.
  2. 2.
  3. 3.
fix $\psi$, update $\theta$ using StaD:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
end while
return $\theta,\psi$.
Algorithm 2 Diff-Instruct++ Pseudo Code.
A.2 Meanings of Hyper-parameters.
@@ A.1 Pseudo-code of Diff-Instruct++
A.3 Experiment Details for Pre-training and Alignment
@@
A.4 More Discussions on Findings of Qualitative Evaluations